# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v1

@@ V-C1 State-based Observations
RFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $0.98$ $\bm{1}$ $\bm{0.96}$ $\bm{1}$ $\bm{1}$ $\bm{0.98}$ $0.96$ $\bm{0.78}$ $\bm{0.82}$ $\bm{0.9}$ $\bm{0.84}$ $\bm{0.9}$ $0.16$ $\bm{0.3}$ $\bm{0.36}$ $\bm{0.56}$ $\bm{0.7}$
SRFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $0.94$ $0.96$ $\bm{1}$ $\bm{0.98}$ $\bm{0.98}$ $0.72$ $0.7$ $0.78$ $0.7$ $0.72$ $\bm{0.26}$ $0.2$ $0.28$ $0.5$ $0.56$
DP 00 $0.78$ $0.96$ $0.96$ $0.98$ 00 $0.38$ $0.82$ $0.92$ $0.9$ 00 $0.4$ $0.62$ $0.66$ $0.66$ 00 00 $0.04$ $0.1$ $0.08$
TABLE VIII: Jerkiness of predicted robot trajectories for different NFE on the robomimic tasks with state-based observations. All values are expressed in thousands, where the lower the smoother the prediction.
Task Lift Can Square Tool Hang
NFE $1$ $2$ $3$ $5$ $10$ $1$ $2$ $3$ $5$ $10$ $1$ $2$ $3$ $5$ $10$ $1$ $2$ $3$ $5$ $10$
RFMP $\bm{9.75}$ $\bm{8.82}$ $9.05$ $\bm{8.82}$ $\bm{8.44}$ $7.92$ $\bm{6.62}$ $6.59$ $\bm{6.45}$ $9.68$ $\bm{9.11}$ $\bm{6.39}$ $\bm{5.49}$ $\bm{5.3}$ $\bm{4.29}$ $\bm{4.43}$ $\bm{3.98}$ $\bm{4.54}$ $\bm{5.25}$ $\bm{4.79}$
SRFMP $10.2$ $10.1$ $\bm{7.79}$ $9.75$ $9.35$ $\bm{7.6}$ $7.75$ $\bm{6.35}$ $7.16$ $\bm{6.86}$ $9.29$ $9.67$ $6.59$ $8.3$ $7.81$ $5.48$ $5.0$ $5.33$ $6.28$ $6.29$
DP $344$ $14.4$ $8.15$ $7.3$ $8.57$ $558$ $19.9$ $6.94$ $5.93$ $6.19$ $558$ $19.9$ $7.28$ $14.7$ $5.48$ $777$ $9.42$ $9.95$ $5.39$ $7.41$
Next, we evaluate the performance of the policies for different NFE in the testing phase. For RFMP and SRFMP, we use the $50$-epoch models for Lift, Can, and Square, and the $100$-epoch model for Tool Hang. DP is further trained for a total of $300$ epochs and we select the model at the best validation epoch. The results are reported in Table VII. RFMP and SRFMP outperform DP for all task and all NFE, even though DP was trained for more epochs. Moreover, we observe that RFMP and SRFMP are generally more robust to low NFE than DP. They achieve $100$% success rate at almost all NFE values for the easier Lift and Can tasks, while DP’s performance drastically drops for $1$ and $2$ NFE. We observe a similar trend for the Square task, where the performance of RFMP and SRFMP slightly improves when increasing the NFE. The performance of all models drops for Tool Hang, which is the most complex of the considered tasks. In this case, the performance of RFMP and SRFMP is limited for low NFE values and improves for higher NFE. DP performs poorly for all considered NFE values. Table VIII reports the jerkiness as a measure of the smoothness of the trajectories generated by the different policies. We observe that RFMP and SRFMP produce arguably smoother trajectories than DP for low NFE, as indicated by the lower jerkiness values. The smoothness of the trajectories becomes comparable for higher NFE. In summary, both RFMP and SRFMP achieve high success rates and smooth action predictions with low NFE, enabling faster inference without compromising task completion.
V-C2 Vision-based Observations
Next, we assess our models performance when the vector field is conditioned on visual observations. We consider the tasks Lift, Can and Square with the same policy settings and networks (see Table I). Each observation ${\bm{o}}^{s}$ at time $s$ corresponds to the embeddings vector obtained from an image of an over-the-shoulder camera and an image of an in-hand camera. We train the models for a total $100$ epochs and use the best-performing checkpoint for evaluation. The performance of different policies is reported in Table IX. RFMP and SRFMP consistently outperform DP on all tasks, regardless of the NFE. As for the previous experiments, our models are remarkably robust to changes in NFE compared to DP. Importantly, SRFMP consistently surpassed RFMP for $1$ and $2$ NFE. Regarding Can and Square tasks, SRFMP with $1$ NFE achieved performance on par with RFMP using $3$ NFE. This efficiency gain showcases the benefits of enhancing the policies with stability to the target distribution for reducing their inference time. We additionally compare RFMP and SRFMP against CP by reporting the performance obtained from [3] in Table IX. Our models achieve a competitive performance compared to CP, which is a method aimed at steeping up inference. However, in contrast to CP, our models are easy and fast to train.
TABLE IX: Success rate as a function of NFE on vision-based robomimic tasks. The CP results correspond to those reported in [3].
Task Lift Can Square
NFE NFE NFE
Policy $1$ $2$ $3$ $5$ $10$ $1$ $2$ $3$ $5$ $10$ $1$ $2$ $3$ $5$ $10$
RFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $0.78$ $0.82$ $\bm{0.9}$ $\bm{0.96}$ $\bm{0.94}$ $0.56$ $0.74$ $\bm{0.9}$ $\bm{0.9}$ $\bm{0.9}$
SRFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{0.88}$ $\bm{0.88}$ $\bm{0.9}$ $0.9$ $0.86$ $\bm{0.86}$ $\bm{0.82}$ $\bm{0.9}$ $0.88$ $\bm{0.9}$
DP 00 $0.7$ $0.96$ $0.98$ $0.98$ 00 $0.38$ $0.66$ $0.68$ $0.66$ 00 $0.04$ $0.16$ $0.26$ $0.12$
CP $\underline{1}$ N.A. $\underline{1}$ N.A. N.A. $\underline{0.98}$ N.A. $\underline{0.95}$ N.A. N.A. $\underline{0.92}$ N.A. $\underline{0.96}$ N.A. N.A.
V-D Real Robotic Experiments
@@
V-D1 Experimental Setup
Figure 7 shows our experimental setup. The tasks are performed on a Franka Emika Panda robot arm. We collect the demonstrations via a teleoperation system consisting of two robot twins. Demonstrations are collected by an expert who guides the source robot. All the demonstrations were recorded in a fairly controlled environment with minimal variation in lighting and background. The target robot then reads the end-effector pose of the source robot and reproduces it via a Cartesian impedance controller. Each observation ${\bm{o}}^{s}$ is composed of the end-effector position, and of the image embedding obtained from the ResNet vision backbone that processes the images from an over-the-shoulder camera. The policies are trained to generate $8$-dimensional actions composed of the position, orientation, and gripper state.
@@ V-D2 Pick & Place
Our experiments also revealed that slight variations in background and lighting conditions (e.g., cloudy and sunny days), had minimal impact on the policies performance. However, consistent failures were observed when the mug was initially positioned on the right side of the yellow mat, where the robot often obstructs the external camera view when approaching the mug. A multi-camera setting may improve performance on such occlusion cases.
V-D3 Mug Flipping
In summary, our findings from both simulated and real-world tasks show that RFMP and SRFMP offer significant advantages over DP. In particular, RFMP and SRFMP achieve faster inference by using fewer NFE without compromising success rates regardless of the observation type. This translates into highly-robust visuomotor policies. Importantly, these advantages do not come at the cost of elaborated training strategies like those used in consistency-based models. In fact, our models demonstrate competitive performance against CP, while being notably easier to train. Regarding the difference between SRFMP and RFMP, the results did not show significant performance gains in terms of success rate and prediction smoothness. Nevertheless, SRFMP shows to be easier to train, achieving higher success rate than RFMP for fewer training epochs (e.g., in Lift, Can, Square tasks).
@@ VI Conclusion
This paper introduced Stable Riemannian Flow Matching Policy (SRFMP), a novel framework that combines the easy training of flow matching with stability-based robustness properties for visuomotor policy learning. SRFMP builds on our extension of stable flow matching to Riemannian manifolds, providing stable convergence of the learned flow to the support of Riemannian target distributions. Our simulated and real-world experiments show that both our previous work on RFMP and its stable counterpart SRFMP outperform diffusion policies and achieve comparable performance to recent distillation-based extensions, while offering advantages in terms of inference speed, ease of training, and robust performance even with limited NFEs and training epochs. Future work will focus on exploring equivariant policy structures to potentially reduce the number of required demonstrations and to improve generalization. Additionally, we aim to investigate multi-modal perception backbones for tackling contact-rich tasks.

### v2

@@ V-C2 Vision-based Observations
RFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $0.78$ $0.82$ $\bm{0.9}$ $\bm{0.96}$ $\bm{0.94}$ $0.56$ $0.74$ $\bm{0.9}$ $\bm{0.9}$ $\bm{0.9}$ $0.6$ $0.82$ $0.78$ $0.78$ $0.8$
SRFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{0.88}$ $\bm{0.88}$ $\bm{0.9}$ $0.9$ $0.86$ $\bm{0.86}$ $\bm{0.82}$ $\bm{0.9}$ $0.88$ $\bm{0.9}$ $0.8$ $0.82$ $0.82$ $0.78$ $0.78$
DP 00 $0.7$ $0.96$ $0.98$ $0.98$ 00 $0.38$ $0.66$ $0.68$ $0.66$ 00 $0.04$ $0.16$ $0.26$ $0.12$ 00 $0.54$ $0.66$ $0.78$ $0.82$
CP $0.5$ $0.44$ $0.46$ $0.46$ $0.4$ $0.82$ $0.82$ $0.84$ $0.84$ $0.82$ $0.32$ $0.52$ $0.54$ $0.48$ $0.52$ $0.4$ $0.36$ $0.44$ $0.38$ $0.36$
V-D Franka Kitchen Benchmark
TABLE XII: Success rate as a function of NFE on the Franka kitchen benchmark.
NFE
Policy 1 2 3 5 10
RFMP $0.04$ $0.08$ $0.12$ $0.14$ $0.14$
SRFMP $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$ $\bm{1}$
DP 00 00 00 $0.02$ $0.02$
We evaluate the capability of RFMP and SRFMP in handling more complex, long-horizon manipulation tasks on the widely-used Franka Kitchen benchmark [58]. We use the dataset from [67], which comprises $19$ expert demonstrations totaling $4209$ time steps and involves a sequential execution of four tasks: open the microwave, put the kettle on the top burner, switch on the light and slide the cabinet. The original dataset considers actions in Euclidean space as joint angular velocities. To evaluate the performance of our Riemannian policy framework, we instead consider end-effector trajectories obtained from joint configurations via forward kinematics. Therefore, each data point is composed of the end-effector position and orientation (as a unit quaternion), and the gripper state. Consequently, the predicted action sequence lies on the product of manifolds $\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{2}$.
Each model is trained for $500$ epochs with a batch size of $32$, and evaluated using the best validation checkpoint over $50$ test episodes with randomized initial states. The performance of different policies is reported in Table XII. SRFMP exhibits consistently high performance across all NFE settings, indicating strong robustness, and significantly outperforms RFMP and DP. The low success rate of RFMP and DP is due to the fact that their trajectories often exhibit local inaccuracies, leading to failing at least one of the four subtasks. This is then reported as a failure for the long-horizon task, leading to low success rates overall. In contrast, the trajectories produced by SRFMP are more precise, leading to the successful completion of all subtasks.
V-E Real Robotic Experiments
@@
V-E1 Experimental Setup
Figure 9 shows our experimental setup. The tasks are performed on a Franka Emika Panda robot arm. We collect the demonstrations via a teleoperation system consisting of two robot twins. Demonstrations are collected by an expert who guides the source robot. The target robot then reads the end-effector pose of the source robot and reproduces it via a Cartesian impedance controller. The demonstration data were smoothed to mitigate the noise arising from the teleoperation setup. Each observation ${\bm{o}}^{s}$ is composed of the end-effector position, and of the image embedding obtained from the ResNet vision backbone that processes the images from an over-the-shoulder camera. The policies are trained to generate $8$-dimensional actions composed of the position, orientation, and gripper state lying on the product of manifold $\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}$.
@@ V-E2 Pick & Place
Notice that the experiments involved natural variations in background and lighting conditions, further exposing the policies to realistic sensor noise. These slight variations in background and lighting conditions (e.g., cloudy and sunny days), had minimal impact on the policies performance. However, consistent failures were observed when the mug was initially positioned on the right side of the yellow mat, where the robot often obstructs the external camera view when approaching the mug. A multi-camera setting may improve performance on such occlusion cases.
@@ V-E3 Mug Flipping
V-F Experiments Summary
We conducted comprehensive evaluations across diverse benchmarks in simulation, namely the Euclidean and Sphere Push-T tasks, the Robomimic benchmark under both state- and vision-based observation, the long-horizon Franka Kitchen benchmark, as well as two real-world manipulation tasks. These tasks display various action spaces with different types of geometric constraints. While the Euclidean Push-T and Robomimic tasks are defined on standard Euclidean spaces — a special type of Riemannian manifolds — the Sphere Push-T task involves an action space constrained to the hypersphere $\mathbb{S}^{2}$, and both the Franka Kitchen and real-world experiments operate on product Euclidean and sphere manifolds, accounting for position, orientation, and gripper state. Through their Riemannian formulation, RFMP and SRFMP are designed to naturally handle Euclidean and non-Euclidean action spaces, thus satisfying constraints such as the unit-norm of quaternions. It is important to emphasize that the Riemannian formulation is necessary not only to handle geometric constraints but also to guarantee stability in SRFMP. Naive approximations, e.g., unit-norm normalization as a post-processing step for quaternions, would break the stability guarantees in SRFMP.
Our findings from both simulated and real-world tasks show that RFMP and SRFMP offer significant advantages over DP. In particular, RFMP and SRFMP achieve faster inference by using fewer NFE without compromising success rates regardless of the observation type. This translates into highly-robust visuomotor policies. Importantly, these advantages do not come at the cost of elaborated training strategies like those used in consistency-based models. In fact, our models outperformed CP, while being notably easier and more efficient to train. Regarding the difference between SRFMP and RFMP, the results did not show significant performance gains in terms of success rate and prediction smoothness, at the exception of the long-horizon task of the Franka kitchen benchmark, where SRFMP significantly outperformed RFMP and DP. Nevertheless, SRFMP shows to be easier to train, achieving higher success rate than RFMP for fewer training epochs (e.g., in Lift, Can, Square tasks).
@@ VI Conclusion
This paper introduced Stable Riemannian Flow Matching Policy (SRFMP), a novel framework that combines the easy training of flow matching with stability-based robustness properties for visuomotor policy learning. SRFMP builds on our extension of stable flow matching to Riemannian manifolds, providing stable convergence of the learned flow to the support of Riemannian target distributions. Our simulated and real-world experiments show that both our previous work on RFMP and its stable counterpart SRFMP outperform diffusion policies and distillation-based extensions, while offering advantages in terms of inference speed, ease of training, and robust performance even with limited NFEs and training epochs. Future work will focus on exploring equivariant policy structures to potentially reduce the number of required demonstrations and to improve generalization. Additionally, we aim to investigate multi-modal perception backbones for tackling contact-rich tasks.